CH/Agen/94/02

Agenda

Charissa Group Meeting

To be held on: Wednesday, 23 Feb 1994, 11.15 am

To be held at:  Birmingham

Distribution: WNC,NMC,RAC,MF,BRF,GJG,SJAH,RAH,CDJ,JK,JSL,ACM, VP,WDMR,GT,RPW,DDW,DLW,JZ, all by e-mail/server

1. Minutes of 5 January 1994
2. Matters Arising / Actions
3. Update on NPC/SERC matters  (DLW)
4. Progress with MEGHA development

1. Prototype Module

2. Chamber Design

3. Detectors

4. Gas Supply

5. Progress with electronics  (WDMR,DLW,RAH)

1. Preamplifiers

2. Amplifiers

3. Power Supplies

6. Progress with Data Acquisition  (RAH,VP,SJAH)

1. ANU running March/April 1994

1. Other Developments

7. Status of plans for ANU running March/April 1994  (DLW)
8. Provisional plans for ANU running August 1994  (BRF)
9. Status of Strasbourg Developments  (JSL)

Refer: Ch/Doc/94/01, Minutes of Strasbourg Meeting

10. Spend on Grants  (DLW,WDMR,WNC,BRF)
11. Detector Stocks  (RPW)
12. SOPHIE and the SHEILAS^{†}  (NMC)
13. Publications and Talks
14. Any other business
15. Date of next meeting

W.N. Catford (Secretary) 22 February 1994

†Don’t blame me for this acronym … Secretary